OpenStudy (anonymous):

A team won 60% of the 20 games it playd so far this season. If the team plays a total of 50 games all season and wins 80% of the remaining games, then what is the % of games it won for the entire season?

OpenStudy (a_clan):

72%

OpenStudy (anonymous):

30 are remaining 80% is 24 60% of 20 is 12 total 36 which is 71. something say that 72

OpenStudy (anonymous):

its exactly 72 x=(36/50)*100 =36*2 =72%